# Sequential Diversification with Provable Guarantees

### v2

@@ 9. Conclusion
Acknowledgements.
We sincerely thank Ilie Sarpe for carefully reviewing our manuscript and providing timely, detailed feedback, which helps to improve the quality of this paper. This research is supported by the ERC Advanced Grant REBOUND (834862), the EC H2020 RIA project SoBigData++ (871042), and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
10. ETHICS STATEMENT
[…]
To ensure the ethical integrity of our work, all datasets utilized in this study are publicly available and have been processed to protect user privacy. Specifically, user IDs have been anonymized and replaced with numeric identifiers, ensuring that no personal or identifying information is disclosed.
@@ Appendix A Omitted proofs from Section 5
See 4
Proof.
We start with demonstrating that the sequential sum-diversity objective can be precisely reformulated via the expression $2\sum_{i=2}^{n}\sum_{j=1}^{i-1}\prod_{t=1}^{i}\mathrm{Pr}_{\pi(t)}d(\pi(i),\pi(j))$. Next, the formulation can be simplified using the definition of $p_{O_{i}}$ and $d(\pi(i),O_{i-1})$.
(A.1) $\sum_{k=1}^{n}\mathrm{Pr}(A=O_{k})\sum_{\pi(i),\pi(j)\in O_{k}}d(\pi(i),\pi(j))$
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{k=1}^{n}\mathrm{Pr}(A=O_{k})\sum_{i,j\leq k}d(\pi(i),\pi(j))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{k=1}^{n-1}(\prod_{t=1}^{k}p_{\pi(t)}-\prod_{t=1}^{k+1}p_{\pi(t)})\sum_{i,j\leq k}d(\pi(i),\pi(j))$
$+\prod_{t=1}^{n}p_{\pi(t)}\sum_{i,j\leq n}d(\pi(i),\pi(j))$
$=\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(-\sum_{i,j\leq k}d(\pi(i),\pi(j))+\sum_{i,j\leq k+1}d(\pi(i),\pi(j)))$
$=\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(\sum_{i=k+1,j\leq k}d(\pi(i),\pi(j))+\sum_{j=k+1,i\leq k}d(\pi(i),\pi(j)))$
$=2\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(\sum_{i=k+1,j\leq k}d(\pi(i),\pi(j)))$
$=2\sum_{k=2}^{n}\prod_{t=1}^{k}p_{\pi(t)}(\sum_{i=k,j\leq k-1}d(\pi(i),\pi(j)))$
$=2\sum_{i=2}^{n}\sum_{j=1}^{i-1}\prod_{t=1}^{i}p_{\pi(t)}d(\pi(i),\pi(j)).$
@@
On the other hand, if there does not exist a size $k$ clique in $G$, then the densest possible graph structure of $G$ contains $T=\left\lfloor\frac{n}{k-1}\right\rfloor$ size-$(k-1)$ cliques, and one clique of size $L=n-(k-1)T$. A node in a size-$(k-1)$ clique can connect to at most $k-2$ nodes in another size-$(k-1)$ clique, thus the maximum number of edges between any two size-$(k-1)$ cliques is $(k-1)(k-2)$. Similarly, the number of edges between a size-$L$ clique and a size-$(k-1)$ clique is $(k-2)L$. Given this graph structure, we can construct an ordered sequence $\tilde{O}=(\tilde{\pi}(i))_{i=1}^{n}$, where sequence $[\tilde{\pi}((j-1)(k-1)+1),\cdots,\tilde{\pi}((k-1)\times j)]$ corresponds to the $j$-th size-$(k-1)$ clique of $G$, with $j\in[1,T]$, and the remaining sequence corresponds to the size-$L$ clique of $G$.
Let $\mathcal{U}(\tilde{O})=\mathcal{S}_{+}(\tilde{O})$. One can verify that $\mathcal{U}(\tilde{O})$ is an upper bound of the optimal objective value of the corresponding MaxSSD problem. It holds that
(A.3) $\mathcal{U}(\tilde{O})$ $=\sum_{i=1}^{n-1}p_{\tilde{O}_{i+1}}d(\tilde{\pi}(i+1),\tilde{O}_{i})$
=$\stackrel{{\scriptstyle(a)}}{{=}}2(p^{2}+2p^{3}+\cdots+(k-2)p^{k-1})$
@@
<$\stackrel{{\scriptstyle(b)}}{{<}}\mathcal{L}(O)+(\epsilon-1)p^{k}+2n(p^{k+1}+p^{k+2}+\cdots+p^{n}),$
where $(a)$ holds because of Lemma 4 and $(b)$ holds because for $j\in[k+1,n]$, the $j$-th node contributes at most $2np^{j}$ to $\mathcal{U}(\tilde{O})$.
@@
$\mathcal{L}(O)-\mathcal{U}(\tilde{O})$ $=\theta-\mathcal{U}(\tilde{O})$
@@
A.1. Proof of Theorem 2
We start by observing that $\mathcal{H}(O)$ can be reformulated according to Lemma 3. The reformulation helps us to relate $\mathcal{H}(O)$ with $\mathcal{S}_{+}(O)$. Notice that $p_{O_{i+1}}$ is the common coefficient of $d_{L}(O_{i+1})$ in $\mathcal{H}(O)$ and $d(\pi(i+1),O_{i})$ in $\mathcal{S}_{+}(O)$. Hence, we essentially only compare $d_{L}(O_{i+1})$ with $d(\pi(i+1),O_{i})$, for all $11\leq i\leq n-$ 1.
Lemma 0.
$\mathcal{H}(O)$ can be equivalently formulated as
$\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}p_{O_{i+1}}d_{L}(O_{i+1}),$
where $d_{L}(O_{k})\coloneqq\sum_{t=1}^{k-1}d(\pi(t),\pi(t+1))$.
The following corollary is a direct implication from the triangle inequality, $2d(\pi(i+1),O_{i})\geq d_{L}(O_{i+1})$, for any sequence $O$.
Corollary 0.
For any ordered sequence $O$ of items in $U$ it holds
$2\mathcal{S}_{+}{(O)}\geq\mathcal{H}(O).$
Let $O^{*}$ be the optimal sequence of $\mathcal{S}_{+}$: $O^{*}=\arg\allowbreak\max\mathcal{S}_{+}{(O)}$. Let $(\pi(i),\pi(i+1))$ be any pair of items swapped in $O^{*}$, and denote $O^{\prime}=(\ldots,\pi(i+1),\pi(i),\ldots)$. By optimality, it follows that $\mathcal{S}_{+}{(O^{*})}\geq\mathcal{S}_{+}{(O^{\prime})}$. With some simplification, we get the following lemma.
Lemma 0.
$d(\pi(i+1),O_{i}^{*})\leq\frac{1-p_{\pi(i+1)}}{p_{\pi(i+1)}}\sum\nolimits_{j=1}^{i}\frac{p_{\pi(j+1)}}{1-p_{\pi(j+1)}}d(\pi(j),\pi(j+1)),$
for any $11\leq i\leq n-$ 1.
We observe that $\frac{1-p_{i}}{p_{i}}\leq\frac{1-a}{a}$, and $\frac{p_{i}}{1-p_{i}}\leq\frac{b}{1-b}$ if $p_{i}\in[a,b]$. The following corollary holds directly by the implication of Lemma 5 that $d(\pi(i+1),O_{i}^{*})\leq\frac{b(1-a)}{a(1-b)}d_{L}(O_{i+1}^{*})$.
Corollary 0.
We are ready to prove Theorem 2.
@@ A.1. Proof of Theorem 2
where inequality ()($\dagger$)( † ) follows from Corollary 6, and ()($\ddagger$)( ‡ ) from Corollary 4. Thus we proved $\frac{\mathcal{S}_{+}{(O^{\alpha})}}{\mathcal{S}_{+}{(O^{*})}}\geq\frac{a(1-b)}{2b(1-a)}\alpha$. ∎
@@
We simply re-arrange the formulas.
(A.4) $\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}w_{i}d(\pi(i),\pi(i+1))$
$=\sum_{i=1}^{n-1}\sum_{k=i+1}^{n}p_{O_{k}}d(\pi(i),\pi(i+1))$
$=p_{O_{2}}d(\pi(1),\pi(2))+p_{O_{3}}(d(\pi(1),\pi(2))+d(\pi(2),\pi(3)))$
$+\cdots+p(O)\sum_{t=1}^{n-1}d(\pi(t),\pi(t+1))$
$=p_{O_{2}}d_{L}(O_{2})+p_{O_{3}}d_{L}(O_{3})+\cdots+p(O)d_{L}(O)$
$=\sum_{i=1}^{n-1}p_{O_{i+1}}d_{L}(O_{i+1}).$
See 4
Proof.
First, we show that if $2d(\pi(i+1),O_{i})\geq d_{L}(O_{i+1})$ holds for any $i\in[n-1]$, then our lemma is proven:
$2\mathcal{S}_{+}(O)$ $=2\sum_{i=1}^{n-1}p_{O_{i+1}}d(\pi(i+1),O_{i})$
$=\sum_{i=1}^{n-1}p_{O_{i+1}}2d(\pi(i+1),O_{i})$
@@
Next, we show that $2d(\pi(i+1),O_{i})\geq d_{L}(O_{i+1})$ holds for any $11\leq i\leq n-$ 1:
$d_{L}(O_{i+1})$ $=\sum_{j=1}^{i}d(\pi(j),\pi(j+1))$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}d(\pi(i+1),O_{i})+d(\pi(i+1),O_{i})-d(\pi(i+1),\pi(1))$
$<2d(\pi(i+1),O_{i}),$
@@
Let $O^{*}=(\pi(i))_{i=1}^{n}$ denote an optimal sequence. To proof this corollary it suffices to show that $d(\pi(i+1),O^{*}_{i})\leq\frac{b(1-a)}{a(1-b)}d_{L}(O^{*}_{i+1})$, for all $i\in[n-1]$.
@@
Substituting Section A.1 into Lemma 5 leads to
$d(\pi(i+1),O_{i}^{*})$ $\leq\frac{1-p_{\pi(i+1)}}{p_{\pi(i+1)}}\sum_{j=1}^{i}\frac{p_{\pi(j+1)}}{1-p_{\pi(j+1)}}d(\pi(j),\pi(j+1))$
@@ Appendix B Additional preliminaries
Let $s_{n}$ be the sum of the the first $n$ terms of a geometric series, namely $s_{n}=\sum_{i=0}^{n-1}ar^{i}$, where $a$ is a constant, and $r$ is any number such $10<r<$ 1. Then, $s_{n}=a\frac{1-r^{n}}{1-r}$.
@@ Appendix C Omitted proofs from Section 6
See 1
Proof.
$\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1)).$
$\mathcal{H}(O^{o})=\sum\nolimits_{i=1}^{n-1}W_{O_{i}^{o}}d(\pi^{o}(i),\pi^{o}(i+1)).$
$\sum_{j=i+1}^{n}p^{j}$ =$\stackrel{{\scriptstyle(a)}}{{=}}p^{i+1}\frac{1-p^{n-i}}{1-p}$
$=\frac{p^{i+1}}{1-p}-\frac{p^{n+1}}{1-p},$
where equality $(a)$ holds by Lemma 1.
Recall that the $\textsc{B}\tau\textsc{I}$ algorithm finds the optimal $\tau$-item sequence that maximizes the following equation:
(C.1) $O_{\tau}=\operatorname*{arg\,max}_{O_{\tau}\sqsubseteq\pi(U)}\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1)).$
For the convenience in our calculations, we let
$\ell_{\pi}=\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1)).$
Let us derive a lower bound on $\mathcal{H}(O)$ as follows:
(C.2) $\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1))-\sum_{i=1}^{n-1}\frac{p^{n+1}}{1-p}d(\pi(i),\pi(i+1))$
$\geq\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1))-\sum_{i=1}^{\tau-1}\frac{p^{n+1}}{1-p}d(\pi(i),\pi(i+1))$
$=\ell_{\pi}-\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1)).$
(C.3) $\mathcal{H}(O^{o})$ $=\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d(\pi^{o}(i),\pi^{o}(i+1))-\sum_{i=1}^{n-1}\frac{p^{n+1}}{1-p}d(\pi^{o}(i),\pi^{o}(i+1))$
$\leq\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d(\pi^{o}(i),\pi^{o}(i+1))$
$\leq\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t\cdot(\tau-1)}\frac{p^{i+1}}{1-p}d(\pi^{o}(i),\pi^{o}(i+1))$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\ell_{\pi}(1+p^{\tau-1}+\cdots+p^{(T-1)(\tau-1)})$
$\leq\frac{\ell_{\pi}}{1-p^{\tau-1}}.$
Combining Equations (C.2) and (C.3), we get
$=(1-p^{\tau-1})\left(1-\frac{\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1))}{\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1))}\right)$
$\geq(1-p^{\tau-1})\left(1-\frac{\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1))}{\sum_{i=1}^{\tau-1}\frac{p^{\tau}}{1-p}d(\pi(i),\pi(i+1))}\right)$
$=(1-p^{\tau-1})(1-p^{n-\tau+1})$
$=1-p^{\tau-1}-p^{n-\tau+1}+p^{n}$
$\geq 1-p^{\tau-1}-p^{n-\tau}+p^{n}.$
We have demonstrated that the $\textsc{B}\tau\textsc{I}$ algorithm provides an asymptotic $(1-p^{\tau-1}-p^{n-\tau}+p^{n})$-approximation. ∎
@@ C.1. Proof of Theorem 2
(C.6) $\begin{split}\mathcal{H}(O)&\stackrel{{\scriptstyle(a)}}{{\geq}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}d(\pi(2i-1),\pi(2i))+W_{O_{2i}}d(\pi(2i),\pi(2i+1))\right)\\ &\stackrel{{\scriptstyle(b)}}{{\geq}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}+\frac{1}{2}W_{O_{2i}}\right)d(\pi(2i-1),\pi(2i))\\ &\stackrel{{\scriptstyle(c)}}{{\geq}}\frac{1}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}+\frac{1}{2}W_{O_{2i}}\right)\sum_{i=1}^{\mathcal{T}_{n}}d(\pi(2i-1),\pi(2i))\\ &\stackrel{{\scriptstyle(d)}}{{\geq}}\frac{p^{t_{n}}}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left((t_{n}-2i+1)+\frac{1}{2}(t_{n}-2i)\right)\sum_{i=1}^{\mathcal{T}_{n}}d(\pi(2i-1),\pi(2i))\\ &=\frac{p^{t_{n}}}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left(-3i+1+\frac{3}{2}t_{n}\right)\sum_{i=1}^{\mathcal{T}_{n}}d(\pi(2i-1),\pi(2i))\\ &=\left(p^{t_{n}}\right)\left(\frac{-3-3\mathcal{T}_{n}}{2}+1+\frac{3}{2}t_{n}\right)\sum_{i=1}^{\mathcal{T}_{n}}d(\pi(2i-1),\pi(2i)).\end{split}$
Inequality $(a)$ holds as we decompose $\mathcal{H}(O)$ into two parts, the series of $W_{O_{2i-1}}d(\pi(2i-1),\pi(2i))$ are related with the top-$\mathcal{T}_{n}$ matching, and the series of $W_{O_{2i}}d(\pi(2i),\pi(2i+1))$ are related with the edges that connect this matching.
@@
$\sum_{i=1}^{\mathcal{T}_{n}}d(\pi(2i-1),\pi(2i))\geq\frac{1}{4}\frac{t_{n}-2}{t_{n}-1}\ell_{t_{n}},$
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{e-1}{e}+\Theta\left(\frac{1}{t_{n}}\right)\right)\left(\frac{1}{e}-\Theta\left(\frac{1}{t_{n}}\right)\right)\left(\frac{3}{16}-\Theta\left(\frac{1}{t_{n}}\right)\right)\left(1-\Theta\left(\frac{1}{t_{n}}\right)\right)$

### v3

@@ Appendix A Omitted Proofs from Section 5
On the other hand, if there does not exist a size $k$ clique in $G$, then the densest possible graph structure of $G$ contains $T=\left\lfloor\frac{n}{k-1}\right\rfloor$ size-(k1)(k-1) cliques, and one clique of size L=n(k1)TL=n-(k-1)T. A node in a size-(k1)(k-1) clique can connect to at most k2k-2 nodes in another size-(k1)(k-1) clique, thus the maximum number of edges between any two size-(k1)(k-1) cliques is (k1)(k2)(k-1)(k-2). Similarly, the number of edges between a size-$L$ clique and a size-(k1)(k-1) clique is (k2)L(k-2)L. Given this graph structure, we can construct an ordered sequence $\tilde{O}=(\tilde{\pi}(i))_{i=1}^{n}$, where sequence $\tilde{\pi}((j-1)(k-1)+1),\cdots,\tilde{\pi}((k-1)\times j)]$ corresponds to the $j$-th size-(k1)(k-1) clique of $G$, with $j\in[1,T]$, and the remaining sequence corresponds to the size-$L$ clique of $G$.
[…]
$\stackrel{{\scriptstyle(a)}}{{=}}2(p^{2}+2p^{3}+\cdots+(k-2)p^{k-1})$
@@
$\stackrel{{\scriptstyle(b)}}{{<}}\mathcal{L}(O)+(\epsilon-1)p^{k}+2n(p^{k+1}+p^{k+2}+\cdots+p^{n}),$
[…]
@@
[…]
@@
See 2
@@
We simply re-arrange the formulas.
(A.3) $\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}w_{i}d\left(\pi(i),\pi(i+1)\right)$
$=\sum_{i=1}^{n-1}\sum_{k=i+1}^{n}p_{O_{k}}d\left(\pi(i),\pi(i+1)\right)$
$=p_{O_{2}}d\left(\pi(1),\pi(2)\right)+p_{O_{3}}(d\left(\pi(1),\pi(2)\right)+d\left(\pi(2),\pi(3)\right))$
$+\cdots+p(O)\sum_{t=1}^{n-1}d\left(\pi(t),\pi(t+1)\right)$
$=p_{O_{2}}d_{L}(O_{2})+p_{O_{3}}d_{L}(O_{3})+\cdots+p(O)d_{L}(O)$
$=\sum_{i=1}^{n-1}p_{O_{i+1}}d_{L}(O_{i+1}).$
@@
First, we show that if $2d\left(\pi(i+1),O_{i}\right)\geq d_{L}(O_{i+1})$ holds for any $i\in[n-1]$, then our lemma is proven:
$2\mathcal{S}_{+}(O)$ $=2\sum_{i=1}^{n-1}p_{O_{i+1}}d\left(\pi(i+1),O_{i}\right)$
$=\sum_{i=1}^{n-1}p_{O_{i+1}}2d\left(\pi(i+1),O_{i}\right)$
@@
Next, we show that $2d\left(\pi(i+1),O_{i}\right)\geq d_{L}(O_{i+1})$ holds for any $1\leq i\leq n-1$:
$d_{L}(O_{i+1})$ $=\sum_{j=1}^{i}d\left(\pi(j),\pi(j+1)\right)$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}d(\pi(i+1),O_{i})+d(\pi(i+1),O_{i})-d(\pi(i+1),\pi(1))$
$<2d\left(\pi(i+1),O_{i}\right),$
@@
Let $O^{*}=(\pi(i))_{i=1}^{n}$ denote an optimal sequence. To proof this corollary it suffices to show that $d\left(\pi(i+1),O^{*}_{i}\right)\leq\frac{b(1-a)}{a(1-b)}d_{L}(O^{*}_{i+1})$, for all $i\in[n-1]$.
@@
Substituting Appendix A into Lemma 4 leads to
$d\left(\pi(i+1),O_{i}^{*}\right)$ $\leq\frac{1-p_{\pi(i+1)}}{p_{\pi(i+1)}}\sum_{j=1}^{i}\frac{p_{\pi(j+1)}}{1-p_{\pi(j+1)}}d\left(\pi(j),\pi(j+1)\right)$
@@ Appendix B Additional Preliminaries
Let $s_{n}$ be the sum of the the first $n$ terms of a geometric series, namely $s_{n}=\sum_{i=0}^{n-1}ar^{i}$, where $a$ is a constant, and $r$ is any number such $0<r<1$. Then, $s_{n}=a\frac{1-r^{n}}{1-r}$.
@@ C.1. Proof of Theorem 2
(C.3) $\begin{split}\mathcal{H}(O)&\stackrel{{\scriptstyle(a)}}{{\geq}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}d\left(\pi(2i-1),\pi(2i)\right)+W_{O_{2i}}d\left(\pi(2i),\pi(2i+1)\right)\right)\\ &\stackrel{{\scriptstyle(b)}}{{\geq}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}+\frac{1}{2}W_{O_{2i}}\right)d\left(\pi(2i-1),\pi(2i)\right)\\ &\stackrel{{\scriptstyle(c)}}{{\geq}}\frac{1}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left(W_{O_{2i-1}}+\frac{1}{2}W_{O_{2i}}\right)\sum_{i=1}^{\mathcal{T}_{n}}d\left(\pi(2i-1),\pi(2i)\right)\\ &\stackrel{{\scriptstyle(d)}}{{\geq}}\frac{p^{t_{n}}}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left((t_{n}-2i+1)+\frac{1}{2}(t_{n}-2i)\right)\sum_{i=1}^{\mathcal{T}_{n}}d\left(\pi(2i-1),\pi(2i)\right)\\ &=\frac{p^{t_{n}}}{\mathcal{T}_{n}}\sum_{i=1}^{\mathcal{T}_{n}}\left(-3i+1+\frac{3}{2}t_{n}\right)\sum_{i=1}^{\mathcal{T}_{n}}d\left(\pi(2i-1),\pi(2i)\right)\\ &=\left(p^{t_{n}}\right)\left(\frac{-3-3\mathcal{T}_{n}}{2}+1+\frac{3}{2}t_{n}\right)\sum_{i=1}^{\mathcal{T}_{n}}d\left(\pi(2i-1),\pi(2i)\right).\end{split}$
Inequality $(a)$ holds as we decompose $\mathcal{H}(O)$ into two parts, the series of $W_{O_{2i-1}}d\left(\pi(2i-1),\pi(2i)\right)$ are related with the top-$\mathcal{T}_{n}$ matching, and the series of $W_{O_{2i}}d\left(\pi(2i),\pi(2i+1)\right)$ are related with the edges that connect this matching.
@@
$\sum_{i=1}^{\mathcal{T}_{n}}d\left(\pi(2i-1),\pi(2i)\right)\geq\frac{1}{4}\frac{t_{n}-2}{t_{n}-1}\ell_{t_{n}},$
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{e-1}{e}+\Theta\left(\frac{1}{t_{n}}\right)\right)\left(\frac{1}{e}-\Theta\left(\frac{1}{t_{n}}\right)\right)\left(\frac{3}{16}-\Theta\left(\frac{1}{t_{n}}\right)\right)\left(1-\Theta\left(\frac{1}{t_{n}}\right)\right)$